OpenStudy (anonymous):

1) When a theater owner charges $5 for admission, 100 people attend. For every $.20 increase in admission, there is a loss of 1 customer. What admission should be charged in order to maximize revenue? (Only a calculus solution with appropriate derivative testing is acceptable) How much is the maximum revenue and how many people attend?

jimthompson5910 (jim_thompson5910):

let x = number of increases in admission

jimthompson5910 (jim_thompson5910):

so if there are no increases in admission, then the cost is y = 5 if there is one increase in admission (of $0.20), then the cost is now y = 5+0.2 = 5.2 if there are two increases in admission(of $0.20), then the cost is now y = 5+2*0.2 = 5.4 etc etc

jimthompson5910 (jim_thompson5910):

in general, the cost y is y = 5 + 0.2x or y = 0.2x + 5 where x is the number of increases in admission

jimthompson5910 (jim_thompson5910):

now if there are 100 people when the price is $5 and 1 person leaves everytime you bump the price up $0.20, then after x price increases you will have 100-x people left

jimthompson5910 (jim_thompson5910):

So the revenue is R = (number of people)*(price) R = (100-x)*(0.2x + 5) R = 100*(0.2x + 5) - x*(0.2x + 5) R = 20x + 500 - 0.2x^2 - 5x R = -0.2x^2 + 15x + 500

jimthompson5910 (jim_thompson5910):

Let me know if this is making sense so far or not

jimthompson5910 (jim_thompson5910):

The revenue function is R(x) = -0.2x^2 + 15x + 500 the derivative of the revenue function is R(x) = -0.2x^2 + 15x + 500 R ' (x) = -0.4x + 15

jimthompson5910 (jim_thompson5910):

Set the revenue function equal to zero and solve for x R ' (x) = -0.4x + 15 0 = -0.4x + 15 -15 = -0.4x -15/(-0.4) = x 37.5 = x x = 37.5 So the max revenue occurs when you increase the price 37.5 times Note: In the real world, you cannot increase the price a fraction of the time, so you would round to the nearest integer to get 37 or 38. So increasing it 37 or 38 times will lead you to the max revenue

OpenStudy (anonymous):

okay so far so good!

jimthompson5910 (jim_thompson5910):

But let's say we go with x = 37.5 When x = 37.5, R(x) is... R(x) = -0.2x^2 + 15x + 500 R(37.5) = -0.2(37.5)^2 + 15(37.5) + 500 R(37.5) = 781.25 This means that theoretically your max revenue is $781.25 However, you can only increase it a whole number of times, so you would either increase it x = 37 or x = 38 times If x = 37, then R(x) = -0.2x^2 + 15x + 500 R(37) = -0.2(37)^2 + 15(37) + 500 R(37) = 781.2 This means that the more realistic max revenue is $781.20

jimthompson5910 (jim_thompson5910):

So if you were making a presentation on this to your boss, then you would recommend that s/he raise the price by $0.20 a total of 37 times and that will lead to a max revenue of $781.20 If your boss raises it 38 times, then the max revenue will still be $781.20...but the revenue will go down after this point. So 38 is the highest you should go

jimthompson5910 (jim_thompson5910):

everything making sense?

OpenStudy (anonymous):

yes it is making sense so far I am just trying to see what the admission price that should be charged is? I see the max revenue just not the price we should charge per ticket

OpenStudy (anonymous):

37*.20+5 for price ticket?

jimthompson5910 (jim_thompson5910):

well if you increase the price 37 times, then you start out at 5 dollars and you add on 37(0.2) dollars ie y = 0.2x + 5 y = 0.2(37) + 5

jimthompson5910 (jim_thompson5910):

exactly you got it

jimthompson5910 (jim_thompson5910):

evaluate that out to find the price per ticket

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

12.60

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

12.40

jimthompson5910 (jim_thompson5910):

off by 20 cents

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (anonymous):

I did 38 the first time!!

jimthompson5910 (jim_thompson5910):

well it turns out that x = 37 and x = 38 give the same R(x) value

jimthompson5910 (jim_thompson5910):

so that explains why you're able to use x = 37 or x = 38 I guess to be fair to the customer (and make the company's image look better), you should go with x = 37

jimthompson5910 (jim_thompson5910):

it's the smarter of the two decisions

jimthompson5910 (jim_thompson5910):

although technically you can still use x = 38 and still get the max revenue just fine

OpenStudy (anonymous):

okay perfect

OpenStudy (anonymous):

So the first two questions are answered in order to get how many people attend all I do is divide the max revenue by the price per ticket?

jimthompson5910 (jim_thompson5910):

or you can plug x = 37 (or x = 38, depending on which x you use) into 100 - x

jimthompson5910 (jim_thompson5910):

remember you start with 100 people when the price is $5

OpenStudy (anonymous):

well that is easier, haha

jimthompson5910 (jim_thompson5910):

then for every increase, 1 person leaves so x increases leaves you with 100 - x people

jimthompson5910 (jim_thompson5910):

either way works

OpenStudy (anonymous):

wonderful, okay ill post the next

jimthompson5910 (jim_thompson5910):

alright